OpenStudy (anonymous):

Backpacks at a sports store are on sale at 30% off the marked price. Tracy bought a backpack that was marked $60. what was the sale price?

OpenStudy (saifoo.khan):

Find it's 30%. Then subtract the result from $60.

OpenStudy (saifoo.khan):

Can you solve it?

OpenStudy (anonymous):

what..

OpenStudy (saifoo.khan):

?

OpenStudy (saifoo.khan):

Are you just looking for the answer?

OpenStudy (anonymous):

no i want to figure it out too so i know.. but i didnt understand what u saiid like what to do first..

OpenStudy (saifoo.khan):

First of all, Find the 30 percent of the total value (which is 60) The result you will get is the "discount price". So Discount price is something which you dont pay. So subtract the result you got from the total value which is 60.

OpenStudy (anonymous):

Original price = sale price + discount = 60 + 30% * 60 = $78

OpenStudy (anonymous):

If the problem asking for the discount price: 60 - 60 * 30% = $42

OpenStudy (anonymous):

thanks:)